OpenStudy (anonymous):

Add and reduce to lowest terms: 3/t + 6/5t

OpenStudy (saifoo.khan):

\[\huge \frac{21}{5t}\]

OpenStudy (anonymous):

To add this, we need to multiply top and bottom of the first term by 5. So we have: 3/t * 5/5 + 6/5t = 15/5t + 6/5t Adding like terms give us: 21/5t Hope this helps. If it does, please choose good answer. http://www.tutorsean.net for free math videos.